Table 1. Basic-Level Curriculum

Bachelor of Science in Computer Science & Engineering

Year & Quarter Course

(Department, Number, Title)

Category (Credit Hours)

Math & Basic Science Engineering Topics Check (√) if Contains Design General Education

Other

Freshman Math 151  5
1^{st} Qtr.  Calculus & Anal. Geo.
Chem 121  5

General Chem.

EG 166  4 (√)

Eng. Graphics & Problem Solving

2^{nd} Qtr. CIS 221 4 (√)

SW Dev. Using Comp.

Math 152  5

Calculus & Anal. Geo.

Chem 125  5

Chem. for Eng.

3^{rd} Qtr. CIS 222 4 (√)

Dev. of SW Comp.

Math 153  5

Calculus & Anal. Geo.

Phys 131  5

Intr. to Physics Year & Quarter Course

(Department, Number, Title)

Category (Credit Hours)

Math & Basic Science Engineering Topics Check (√) if Contains Design General Education

Other

Sophomore CIS 321  4 (√)
1^{st} Qtr.  Design & Analysis of

Comp. SW

Math 254  5

Calculus & Anal. Geo.

Math 366  3

Discrete Math

Phys 132  5

Intr. to Physics

2^{nd} Qtr. CIS 360 4 (√)

Int. Comp. Sys.

Math 415  4

Diff. Equations

MSE 405  3

Int. Mat. Sc. & Eng.

Phys 133  5

Intr. to Physics

3^{rd} Qtr. CIS 570 3 (√)

File Analysis

EM 220  5

Eng. Mechanics

EE 261  3 (√)

Int. to Logic Des

CIS 459.XX  1

Various Lanuages Year & Quarter Course

(Department, Number, Title)

Category (Credit Hours)

Math & Basic Science Engineering Topics Check (√) if Contains Design General Education

Other

Junior  CIS 560  5 (√)
1^{st} Qtr.  System SW Des.
CIS 675.01  3

Computer Arch.

Stat 427  3

Intr. to Prob. & Stat.

EE 206  1 (√)

Switch. Circ. Lab.

2^{nd} Qtr. CIS 660 3

Int. to Oper. Sys.

CIS 625  3

Automata & Lang.

Stat 428  3

Intr. to Prob. & Stat.

EE 309  1 (√)

Elec.. Circ. Lab.

EE 300  3

Elec. Circuits

3^{rd} Qtr. CIS 541 3

Numerical Methods

CIS 680  3

Data Structures

EE 320  3

Electronic Dev.

ME 500  3

Eng Thermal Sc.

Year & Quarter Course

(Department, Number, Title)

Category (Credit Hours)

Math & Basic Science Engineering Topics Check (√) if Contains Design General Education

Other

Senior  CIS 655  4 (√)
1^{st} Qtr.  Prog. Languages
EE 567  2 (√)

Microprocessor Lab.

ISE 504  3

Eng. Economic Analy.

Technical Elective

2^{nd} Qtr.  CIS 601  1

Social & Ethical Issues in Comptng.

Technical Elective

3^{rd} Qtr.  Capstone Des.  3 or 4 (√)

Technical Elective The following courses tend to be distributed over the four years with different students taking these courses at different times:

Year & Quarter Course

(Department, Number, Title)

Category (Credit Hours)

Math & Basic Science Engineering Topics Check (√) if Contains Design General Education

Other

Varies.  Eng 110  5

First Year English Composition

. Second Writing 5

Course

Social Sciences  9
Historical Survey  10
Arts & Humanities  9
Addnl. Gen. Edu.  7
Technical Elective  18